# HumanSense: From Multimodal Perception to Empathetic Context-Aware Responses through Reasoning MLLMs

### v2

@@ Abstract
While Multimodal Large Language Models (MLLMs) show immense promise for achieving truly human-like interactions, progress is hindered by the lack of fine-grained evaluation frameworks for human-centered scenarios, encompassing both the understanding of complex human intentions and the provision of empathetic, context-aware responses. Here we introduce HumanSense, a comprehensive benchmark designed to evaluate the human-centered perception and interaction capabilities of MLLMs, with a particular focus on deep understanding of extended multimodal contexts and the formulation of rational feedback. Our evaluation reveals that leading MLLMs still have considerable room for improvement, particularly for advanced interaction-oriented tasks. Supplementing visual input with audio and text information yields substantial improvements, and Omni-modal models show advantages on these tasks. Furthermore, we argue that appropriate feedback stems from a contextual analysis of the interlocutor’s needs and emotions, with reasoning ability serving as the key to unlocking it. Accordingly, we employ a multi-stage, modality-progressive reinforcement learning to enhance the reasoning abilities of an Omni model, achieving substantial gains on evaluation results. Additionally, we observe that successful reasoning processes exhibit highly consistent thought patterns. By designing corresponding prompts, we also enhance the performance of non-reasoning models in a training-free manner. Project page: https://github.com/antgroup/HumanSense
@@ Introduction
Multimodal Large Language Models (MLLMs) (Xu et al. 2025; Hurst et al. 2024; Anthropic 2024; Team et al. 2023) represent a promising pathway toward realizing this vision. Their ability to holistically process visual, auditory, and textual information enables a comprehensive understanding of users and environments. MLLMs also have the potential to deeply analyze perceived information (Guo et al. 2025) and subsequently plan appropriate feedback, which is not limited to textual responses, but can include suitable emotions, tones, and gesture labels in temporal sequences. Such outputs can be further integrated with video generation (Meng et al. 2025; Qin et al. 2025), speech synthesis (Wang et al. 2023; Suno AI 2023), and talking head (Li et al. 2024c, b; Zhang et al. 2024c; Zheng et al. 2021) methods to provide a highly anthropomorphic interactive experience.
@@
Figure 1: Demonstration of tasks in HumanSense. Note: The questions above are slightly simplified for clarity and brevity.
@@
  • Familiarity Recognition (FR1) evaluates the model’s ability to understand human interpersonal interactions and perceive interpersonal closeness. Based on the duration of acquaintance between conversational participants in The Skin Deep channel, we construct single-choice questions to have the model determine the familiarity level between individuals in the videos.
  • Rapport Recognition (RR1) evaluates the model’s ability to perceive whether the communication atmosphere is harmonious. We annotate video chat content from The Skin Deep channel across multiple dimensions, including interaction frequency, communication atmosphere, and degree of viewpoint conflict, to construct single-choice questions. The evaluation encompasses whether the interactive atmosphere is pleasant and whether conflicts exist in the conversational content.
  • Relation Recognition (RR2) evaluates the model’s ability to determine human relationships through multi-modal context. Based on the relationships between conversational participants in The Skin Deep channel, we construct single-choice questions to have the model predict the relationship type between individuals in the videos, such as “Married”,“Siblings”,“Son/Mother”, etc.. The model must integrate information from visual appearance, age differences, and dialogue content to make a judgment.
@@
  • Fraud Recognition (FR2) evaluates the detection of specific fraudulent behaviors through approximately one-minute phone recordings, derived from the Telecom Fraud Texts dataset (Li, Zhang, and Jiang 2024). The fraud types include “Loan and credit card agency fraud”, “Impersonation of public security, judiciary, and government agencies” and “Impersonation of leaders or acquaintances”, etc.
@@
  • Psychological Chat (PC) evaluates the ability of MLLMs to generate appropriate responses in complex, long-context interactions. Here we use a professional online psychological dataset, Emotional First Aid dataset (Hai Liang Wang 2020), from which we constructed single-choice questions. The model is expected to select an appropriate response based on the previous multi-turn dialogue.
@@
Human Level 0.875 0.917 0.933 0.767 0.933 0.967 0.967 0.889 0.900 0.900 0.800 0.533 0.967 0.833 0.867 0.933
GPT-4o 0.552 0.583 0.233 0.700 0.517 0.733 0.767 0.522 0.400 0.833 - 0.300 0.467 0.467 0.667 -
InternVL3-8B 0.558 0.417 0.467 0.533 0.433 0.833 0.767 0.567 0.333 0.733 - 0.667 0.433 0.433 0.633 -
Qwen2.5-Omni-7B 0.578 0.467 0.500 0.300 0.383 0.633 0.800 0.467 0.600 0.733 0.700 0.567 0.767 0.600 0.700 0.367
Qwen2-Audio-7B - - - - - - - - - - 0.333 - - - - 0.333
@@
LLaVA-Next-Video-7B 0.479 0.500 0.480 0.263 0.583 0.440 0.413 0.264 0.487 0.665 - 0.505 0.560 0.500 0.577 -
Qwen2-VL-7B 0.507 0.473 0.470 0.307 0.322 0.600 0.591 0.424 0.537 0.665 - 0.515 0.627 0.495 0.570 -
Qwen2.5-VL-7B 0.512 0.540 0.497 0.267 0.448 0.550 0.644 0.461 0.523 0.480 - 0.545 0.627 0.485 0.590 -
VideoLLaMA3-7B 0.520 0.543 0.463 0.323 0.517 0.530 0.694 0.561 0.493 0.610 - 0.515 0.587 0.400 0.530 -
LLaVA-OneVision-7B 0.521 0.545 0.510 0.400 0.592 0.620 0.676 0.268 0.503 0.600 - 0.530 0.560 0.430 0.543 -
InternVL3-8B 0.561 0.393 0.483 0.387 0.547 0.670 0.751 0.630 0.567 0.735 - 0.555 0.527 0.490 0.557 -
@@
Ola-7B 0.525 0.557 0.463 0.263 0.573 0.320 0.420 0.371 0.597 0.785 0.733 0.565 0.653 0.640 0.577 0.567
IXC2.5-OmniLive-7B 0.544 0.508 0.467 0.257 0.338 0.660 0.584 0.544 0.533 0.780 0.415 0.500 0.560 0.470 0.463 0.324
Qwen2.5-Omni-7B 0.554 0.473 0.513 0.303 0.350 0.600 0.630 0.438 0.600 0.770 0.740 0.550 0.713 0.650 0.607 0.399
Table 1: Evaluation on HumanSense. Indicates results on the HumanSense (tiny) set, for comparison with human-level performance. While GPT-4o is designed as an omni-modal, the absence of audio input support in its current API precluded its evaluation on audio-related tasks (FR2, PC). For a fair comparison across all models, the overall average scores (Avg.) for omni-modals are calculated without the results from the two audio tasks. The best results for each metric are bolded.
@@
We conduct a comprehensive evaluation of leading Multimodal Large Language Models (MLLMs) with sizes up to 10B, including: (1) Visual LLMs, which represent the most mainstream branch of MLLMs today; (2) Audio LLMs; and (3) Omni-modal LLMs that are natively designed for integrating vision, audio, and text. For Visual LLMs, we assess models such as Qwen2.5-VL (Team 2025), Qwen2-VL (Wang et al. 2024a), LLaVA-OneVision (Li et al. 2024a), LLaVANeXT-Video (Liu et al. 2024), VideoLLaMA3 (Zhang et al. 2025) and InternVL3 (Zhu et al. 2025). In the omni-modal models, we test Qwen2.5-Omni (Xu et al. 2025), IXC2.5OmniLive (Zhang et al. 2024b) and Ola (Liu et al. 2025). For audio LLMs, we evaluate Qwen2-Audio (Chu et al. 2023). Additionally, we test the powerful omni-model GPT-4o (Hurst et al. 2024). All evaluations are conducted in a zero-shot setting, employing the default prompts provided for each model. For video processing, we adhere to the respective official configurations for key parameters, including frame extraction methods, frames per second (FPS), and resolution. We will release our data and code to the community, aiming to facilitate broader evaluation across a diverse range of models.
@@ Results
Specifically, in the first stage, we train using pure video frames and textual Question-Answer(QA) pairs to establish reasoning capabilities that integrate visual perception. In the second stage, we train with audio-based QA to develop reasoning that incorporates auditory perception. Finally, we utilize complete video-audio QA to reinforce reasoning that combines both visual and auditory perception. More training details are included in the supplementary materials.
As shown in Table 2, for vision-centric L1 tasks, stage-1 training already yields significant improvements, indicating that reasoning grounded in visual perception can enhance performance on such tasks. For the audio-related tasks PC and FR2, stage-2 training leads to notable gains over stage-1, revealing the success of incorporating auditory reasoning. Most tasks achieve optimal performance after completing all 3 training stages. We sample correctly answered examples and find that the model is indeed capable of deep thinking by integrating characteristics, emotions, and contextual information, as shown in Figure 5.

### v3

@@ Abstract
While Multimodal Large Language Models (MLLMs) show immense promise for achieving truly human-like interactions, progress is hindered by the lack of fine-grained evaluation frameworks for human-centered scenarios, encompassing both the understanding of complex human intentions and the provision of empathetic, context-aware responses. Here we introduce HumanSense, a comprehensive benchmark designed to evaluate the human-centered perception and interaction capabilities of MLLMs, with a particular focus on deep understanding of extended multimodal contexts and the formulation of rational feedback. Our evaluation reveals that leading MLLMs still have considerable room for improvement, particularly for advanced interaction-oriented tasks. Supplementing visual input with audio and text information yields substantial improvements, and Omni-modal models show advantages on these tasks. Furthermore, grounded in the observation that appropriate feedback stems from a contextual analysis of the interlocutor’s needs and emotions, we posit that reasoning ability serves as the key to unlocking it. We devise a multi-stage, modality-progressive reinforcement learning approach, resulting in HumanSense-Omni-Reasoning, which substantially enhances performance on higher-level understanding and interactive tasks. Additionally, we observe that successful reasoning processes appear to exhibit consistent thought patterns. By designing corresponding prompts, we also enhance the performance of non-reasoning models in a training-free manner. Project page: https://digital-avatar.github.io/ai/HumanSense/
@@ Introduction
Multimodal Large Language Models (MLLMs) (Xu et al. 2025; Hurst et al. 2024; Anthropic 2024; Team et al. 2023) represent a promising pathway toward realizing this vision. Their ability to holistically process visual, auditory, and textual information enables a comprehensive understanding of users and environments. MLLMs also have the potential to deeply analyze perceived information (Guo et al. 2025) and subsequently plan appropriate feedback, which is not limited to textual responses, but can include suitable emotions, tones, and gesture labels in temporal sequences. Such outputs can be further integrated with video generation (Meng et al. 2025; Qin et al. 2025b), speech synthesis (Wang et al. 2023; Suno AI 2023), and talking head (Li et al. 2024c, b; Zhang et al. 2024c; Zheng et al. 2021) methods to provide a highly anthropomorphic interactive experience.
@@
Figure 1: HumanSense benchmark is structured hierarchically to evaluate the end-to-end, human-centric process of perception, understanding, and feedback through a series of demonstrated tasks and questions.
@@
  • Familiarity Recognition (FR) evaluates the model’s ability to understand human interpersonal interactions and perceive interpersonal closeness. Based on the duration of acquaintance between conversational participants in The Skin Deep channel, we construct single-choice questions to have the model determine the familiarity level between individuals in the videos.
  • Rapport Recognition (RR) evaluates the model’s ability to perceive whether the communication atmosphere is harmonious. We annotate video chat content from The Skin Deep channel across multiple dimensions, including interaction frequency, communication atmosphere, and degree of viewpoint conflict, to construct single-choice questions. The evaluation encompasses whether the interactive atmosphere is pleasant and whether conflicts exist in the conversational content.
  • Relation Recognition (RG) evaluates the model’s ability to determine human relationships through multi-modal context. Based on the relationships between conversational participants in The Skin Deep channel, we construct single-choice questions to have the model predict the relationship type between individuals in the videos, such as “Married”,“Siblings”,“Son/Mother”, etc.. The model must integrate information from visual appearance, age differences, and dialogue content to make a judgment.
@@
  • Fraud Recognition (FG) evaluates the detection of specific fraudulent behaviors through approximately one-minute phone recordings, derived from the Telecom Fraud Texts dataset (Li, Zhang, and Jiang 2024). The fraud types include “Loan and credit card agency fraud”, “Impersonation of public security, judiciary, and government agencies” and “Impersonation of leaders or acquaintances”, etc.
@@
  • Psychological Chat (PC) evaluates the ability of MLLMs to generate appropriate responses in complex, long-context interactions. Here we use a professional online psychological dataset, Emotional First Aid dataset (Wang, Wu, and Lang 2020), from which we constructed single-choice questions. The model is expected to select an appropriate response based on the previous multi-turn dialogue.
@@
Human Level 0.875 0.874 0.917 0.933 0.767 0.933 0.967 0.967 0.889 0.900 0.900 0.800 0.533 0.967 0.833 0.867 0.933
GPT-4o 0.552 - 0.583 0.233 0.700 0.517 0.733 0.767 0.522 0.400 0.833 - 0.300 0.467 0.467 0.667 -
InternVL3-8B 0.558 - 0.417 0.467 0.533 0.433 0.833 0.767 0.567 0.333 0.733 - 0.667 0.433 0.433 0.633 -
Qwen2.5-Omni-7B 0.578 0.572 0.467 0.500 0.300 0.383 0.633 0.800 0.467 0.600 0.733 0.700 0.567 0.767 0.600 0.700 0.367
Qwen2-Audio-7B - - - - - - - - - - - 0.333 - - - - 0.333
@@
LLaVA-Next-Video-7B 0.479 - 0.500 0.480 0.263 0.583 0.440 0.413 0.264 0.487 0.665 - 0.505 0.560 0.500 0.577 -
Qwen2-VL-7B 0.507 - 0.473 0.470 0.307 0.322 0.600 0.591 0.424 0.537 0.665 - 0.515 0.627 0.495 0.570 -
Qwen2.5-VL-7B 0.512 - 0.540 0.497 0.267 0.448 0.550 0.644 0.461 0.523 0.480 - 0.545 0.627 0.485 0.590 -
VideoLLaMA3-7B 0.520 - 0.543 0.463 0.323 0.517 0.530 0.694 0.561 0.493 0.610 - 0.515 0.587 0.400 0.530 -
LLaVA-OneVision-7B 0.521 - 0.545 0.510 0.400 0.592 0.620 0.676 0.268 0.503 0.600 - 0.530 0.560 0.430 0.543 -
InternVL3-8B 0.561 - 0.393 0.483 0.387 0.547 0.670 0.751 0.630 0.567 0.735 - 0.555 0.527 0.490 0.557 -
@@
Ola-7B 0.525 0.539 0.557 0.463 0.263 0.573 0.320 0.420 0.371 0.597 0.785 0.733 0.565 0.653 0.640 0.577 0.567
IXC2.5-OmniLive-7B 0.544 0.494 0.508 0.467 0.257 0.338 0.660 0.584 0.544 0.533 0.780 0.415 0.500 0.560 0.470 0.463 0.324
Qwen2.5-Omni-7B 0.554 0.559 0.473 0.513 0.303 0.350 0.600 0.630 0.438 0.600 0.770 0.740 0.550 0.713 0.650 0.607 0.399
Table 1: Evaluation on HumanSense. Indicates results on the HumanSense (tiny) set, for comparison with human-level performance. While GPT-4o is designed as an omni-modal, the absence of audio input support in its current API precluded its evaluation on audio-related tasks (FG, PC). We report two overall average scores: Avg., which excludes the two audio tasks to ensure a fair cross-model comparison, and Avg.*, which is the average across all tasks.
@@
We conduct a comprehensive evaluation of leading Multimodal Large Language Models (MLLMs) with sizes up to 10B, including: (1) Visual LLMs, which represent the most mainstream branch of MLLMs today; (2) Audio LLMs; and (3) Omni-modal LLMs that are natively designed for integrating vision, audio, and text. For Visual LLMs, we assess models such as Qwen2.5-VL (Team 2025), Qwen2-VL (Wang et al. 2024a), LLaVA-OneVision (Li et al. 2024a), LLaVANeXT-Video (Liu et al. 2024), VideoLLaMA3 (Zhang et al. 2025) and InternVL3 (Zhu et al. 2025). In the omni-modal models, we test Qwen2.5-Omni (Xu et al. 2025), IXC2.5OmniLive (Zhang et al. 2024b) and Ola (Liu et al. 2025). For audio LLMs, we evaluate Qwen2-Audio (Chu et al. 2023). Additionally, we test the powerful omni-model GPT-4o (Hurst et al. 2024). All evaluations are conducted in a zero-shot setting, employing the default prompts provided for each model. For video processing, we adhere to the respective official configurations for key parameters, including frame extraction methods, frames per second (FPS), and resolution. We release our data and code to the community, aiming to facilitate broader evaluation across a diverse range of models.
@@ Results
Specifically, in the first stage, we train using pure video frames and textual Question-Answer(QA) pairs to establish reasoning capabilities that integrate visual perception. In the second stage, we train with audio-based QA to develop reasoning that incorporates auditory perception. Finally, we utilize complete video-audio QA to reinforce reasoning that combines both visual and auditory perception. Further details, such as reward functions, data sizes, and key hyperparameters, are provided in the Appendix. The complete training configurations are available in our released code.
As shown in Table 2, for vision-centric L1 tasks, stage-1 training already yields significant improvements, indicating that reasoning grounded in visual perception can enhance performance on such tasks. For the audio-related tasks PC and FG, stage-2 training leads to notable gains over stage-1, revealing the success of incorporating auditory reasoning. Most tasks achieve optimal performance after completing all 3 training stages. We sample correctly answered examples and find that the model is indeed capable of deep thinking by integrating characteristics, emotions, and contextual information, as shown in Figure 5.
@@ Conclusion
Appendix
TRAINING DETAILS
All three training stages are conducted on 8 H20 GPUs. Two types of rewards are employed: Format Reward and Accuracy Reward.
  • Format Reward incentivizes the model to generate predefined structural and labels. It gives a reward of 0 or 1 based on whether the output contains the complete tags of <think></think>. and <answer></answer>.
  • Accuracy Reward evaluates the correctness of the outputs. It provides direct feedback based on how well the model’s predictions align with ground truth or expected results.
Both rewards are assigned equal weight. The learning rate is set to $1\times 10^{-6}$, and the gradient accumulation step is set to 1.
The details of the three training stages are as follows:
  • Stage 1: 2,663 video-only samples (without audio) are used for training over 400 steps.
  • Stage 2: 1,000 audio-only samples are used for training over 150 steps.
  • Stage 3: 1,283 multimodal samples (video and audio) are used for training over 200 steps.
Detailed scores for each task are presented in Table 3.
EVALUATION PROTOCALS
During large model evaluation, besides video or audio inputs, we employ the following prompt to guide the model in answering multiple-choice questions:
And we use the following prompt to have the model perform one-choice question:
And we use the following prompt to have the model perform open-ended question (Human counting with Reid):
For the trained model, we augment each question with a prompt designed to stimulate its reasoning process, which is as follows:
MORE DETAILS OF DATA CONSTRUCTION
For tasks such as Person Appearance, Action Recognition, Gesture Recognition, Fraud Recognition, Emotion Recognition, Abnormal Classification, and Lie Detection, we construct multiple-choice questions based on the built-in annotations from the original dataset (CelebV-HQ, AVA, Jester, Telecom Fraud Texts, CelebV-HQ, ActivityNet, UCF-Crime100, and SEUMLD).
For a single question, the video clip’s corresponding ground truth (GT) label serves as the correct option, while distractor options are randomly sampled from the remaining choices in the answer set. A multiple-choice question is then constructed by combining the video with the correct option and the sampled distractors.
Human Counting with ReID.
We generate question-answer (QA) pairs using the multi-object tracking dataset TAO. We filter the TAO videos to exclude those without human subjects, then segment the remaining videos into 30-50 second clips.
For each segment, we identify the corresponding annotations from the original labels and calculate the total number of unique individuals present. This count serves as the numerical answer for our designed task, forming a complete question-answer pair.
Talking Person Recognition.
We construct questions based on videos from RealTalk. We filter the clips by discarding segments where only one person speaks continuously. For each selected video that contains a dialogue, we extract a single quote. This quote is then used to pose the question: “Who said this?” The objective of this task is to test a large model’s multimodal comprehension ability, specifically its capacity to synchronize audio with visual information to identify the speaker.
Relation Recognition.
The show ”The Skin Deep” displays title cards at the beginning of each episode, indicating the relationship between the two speakers. To create our question-answer pairs, we process the introductory segment of each video frame-by-frame using Optical Character Recognition (OCR) to extract this relationship information, serving as the ground truth for the question: “What is the relationship between the two people in the video?”
Familiarity Recognition.
At the beginning of “The Skin Deep” episodes, on-screen text indicates the duration of the two speakers’ acquaintance. This text serves as the ground truth for the question: ”What is the level of familiarity between the two individuals in the video?”
Rapport Recognition.
We segment clips from “The Skin Deep” into 100-150 second lengths, then conduct a two-level assessment of each segment using both the video and its corresponding transcripts. The first level of video assessment evaluates the conversational interaction based on the following criteria:
  • Interactivity: The dialogue should feature frequent back-and-forth between participants. Points are deducted for long, awkward silences or extended pauses for thought.
  • Emotion: Positive and shared emotions, such as happiness, humor, and relaxation—along with clear emotional resonance between the speakers are scored positively. Divergent or mismatched emotional states between participants result in a lower score.
  • Balance: Contributions to the conversation should be roughly equal from both parties. Scenarios where one individual dominates the conversation while the other is primarily a passive listener are penalized.
  • Coherence: The discussion should remain focused on a consistent topic. Frequent and abrupt switching between topics will negatively impact the score.
Furthermore, we employ large language models to directly rate the harmony of the ASR transcripts, identifying any disagreements.
Emo Strategy.
We use pre-segmented two-person conversation clips from the RealTalk dataset, which are originally used for training speaker generation models. For our task, we select the “talker” from each clip as the input. The corresponding ground truth label is the “listener”s’ facial expression, which we determine using a facial expression recognition algorithm. We refer to this as our ’emotion strategy’ target.
Psychological Chat.
We first filter high-quality advice and responses from online dialogue data of psychologists. This is done using the following prompt:
For the selected responses, we trace back more than 10 dialogue turns to include in the question, aiming to provide sufficient contextual information. Meanwhile, we reference the high-quality responses and generate distractor options.
Models Avg. AR ER GR PA AC CM HC TR FR1 FR2 LD RR1 RR2 ES PC
Baseline 0.559 0.473 0.513 0.303 0.350 0.600 0.630 0.438 0.600 0.770 0.740 0.550 0.713 0.650 0.607 0.399
First Stage (v) 0.584 0.693 0.493 0.427 0.608 0.530 0.626 0.438 0.597 0.720 0.557 0.540 0.707 0.620 0.593 0.540
Second Stage (a) 0.601 0.700 0.510 0.407 0.600 0.550 0.630 0.492 0.587 0.775 0.687 0.545 0.693 0.625 0.593 0.625
Third Stage (v+a) 0.619 0.705 0.513 0.420 0.615 0.550 0.633 0.488 0.740 0.780 0.687 0.555 0.673 0.690 0.620 0.619
Table 3: Evaluation of Optimization Results based on Qwen2.5-Omni-7B. We provide the scores for each task.
Supplement the section on MLLM benchmarks in related works
With the development of multimodal large models, their influence has permeated various fields (Qin et al. 2025a; Wang et al. 2025b, 2024d, a; Li et al. 2025; Wang et al. 2022), leading to the emergence of numerous evaluation benchmarks (Chen et al. 2024; Fu et al. 2025a; Wang et al. 2024c; Li et al. 2024d; Zhang et al. 2024a) have emerged. While they have been successful in general visual understanding, they rarely focus on the subtleties of human perception, comprehension, and planning. In the field of human behavior analysis, there are some specialized datasets. For example, in emotion recognition, AffectNet (Mollahosseini, Hasani, and Mahoor 2017) and in action recognition, Kinetics (Smaira et al. 2020). While these datasets are crucial, they are typically designed for specific classification or detection tasks and do not provide a framework for evaluating a model’s comprehensive reasoning ability. HumanOmniV2 (Yang et al. 2025) focuses on deciphering intentions, interpreting emotions, and detecting potential deception from video. In summary, existing benchmarks either lack multimodal information, focus on general scenarios rather than human-centric ones, or are too task-specific. To systematically evaluate the capabilities of large models in complex human-centered scenarios, we have developed HumanSense.